OpenStudy (anonymous):

someone what am i doing wrong? Joe has a collection of nickels and dimes that is worth $6.10. If the number of dimes was tripled and the number of nickels was decreased by 49, the value of the coins would be $14.05. How many nickels and dimes does he have? my answe.1X+.05y=6.1 3(.1x)+.05y-49(.05)=14.05 .3x+.05y-2.45=14.05 .3x+.05y-.2.45-2.45=14.05+2.45 .3x+.05y-.05y=16.50-.05y .3x/.3=16.5/.3-.05y/.3 x=55+.16666667y

OpenStudy (anonymous):

you aren't doing anything wrong. solve the two equations.

OpenStudy (m):

plug in x into initial equation and solve for y then for x

OpenStudy (anonymous):

yah but the last equation is endless

OpenStudy (anonymous):

.166666666666666666666666666666666666666666666666666666666666666666666666.....

OpenStudy (anonymous):

why dont you try elimination of variable instead of substitution.

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

0.1X+0.05y=6.1 --------------------equation 1 0.3x+0.05y=16.50 -----------------equation 2 multiply equation 1 by 3 and subtract equation 2 from it.

OpenStudy (anonymous):

ooo like a matrix?

OpenStudy (anonymous):

errrrr this doesnt make sense either

OpenStudy (anonymous):

0.1X+0.05y=6.1 --------------------equation 1 0.3x+0.05y=16.50

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

oooo it didnt save my work

OpenStudy (anonymous):

multiply equation 1 by 3.

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

and post what you got.

OpenStudy (anonymous):

i have to redo it now cause it didnt save my work to show you

OpenStudy (anonymous):

0.3x+0.15y = 18.3 is what you get if you multiply equation 1 by 3. so lets call 0.3x+0.15y = 18.3 ------- equation 3. subtract equation 2 from equation 3 and post what you got.

OpenStudy (anonymous):

-.1(.1x+.05y)=-.1(6.1) .3x+0.05y=16.50 -.1x-.05y=-.61 .3x+.05y=16.5 -.04x=-17.11 -.04x/-.04=-17.11/-.04 x=427.75

OpenStudy (anonymous):

this crap just doesnt make sense....

OpenStudy (anonymous):

why did you multiply the equation by 0.1?

OpenStudy (anonymous):

just subtract equation 1 from equation 2 and post what you got. use by equation notations.

OpenStudy (anonymous):

i meant use my equation notation.

OpenStudy (anonymous):

im more confused then when i started...

OpenStudy (anonymous):

0.1X+0.05y=6.1 --------------------equation 1 0.3x+0.05y=16.50 -----------------equation 2 subtract equation 1 from equation 2. post what you get.

OpenStudy (anonymous):

let me do it for you. equation 2 - equation 1 (0.3x+0.05y)-(0.1X+0.05y) = 16.50 - 6.1 make sense?

OpenStudy (anonymous):

.2x+.1y=-10.4

OpenStudy (anonymous):

oh wait .2x=-10.4

OpenStudy (anonymous):

dude, you gotta work with me here. you subtracted 0.l3x - 0.1 x to get 0.2x. why did you ADD 0.05 y and 0.05y? you subtract everything, not just the variable you like. you do 0.05y-0.05y to get 0

OpenStudy (anonymous):

come on man! 0.2x = +10.4, not -10.4 16.5-6.1 = 10.4 why did you get -10.4?

OpenStudy (anonymous):

6.1-16.5

OpenStudy (anonymous):

no, you are subtracting equation 1 FROM equation 2. that means you subtract 6.1 FROM 16.5

OpenStudy (anonymous):

ok fine 10.4

OpenStudy (anonymous):

right, so 0.2x = 10.4 which means x = 10.4/0.2 = 104/2 = 52

OpenStudy (anonymous):

rather, x = 52 in equation 1

OpenStudy (anonymous):

1(54)+.05y=6.1 54+.05y=6.1 54-54+.05y=6.1-54 .05y=-47.9

OpenStudy (anonymous):

oops .05y/.05=-47.9/.05 y=-958

OpenStudy (anonymous):

0.1X+0.05y=6.1 substitute x = 52 in equation 1 you get 0.1(52) + 0.05y = 6.1 5.2+0.05y = 6.1 0.05y = 6.1-5.2 0.05y=0.9 y = 0.9/0.05 y = 90/5 = 18 how can you get y = -958? how can you have negative nine hundred and fifty eight nickels? think!

OpenStudy (anonymous):

oh dear lord my calculator sucks

OpenStudy (anonymous):

you don't need a calculator for this. you multiplied 54 by 1 instead of multiplying it by 0.1

OpenStudy (anonymous):

0

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

no i meant oh

OpenStudy (anonymous):

k so plug 18 in the first or second equation?

OpenStudy (anonymous):

no you got your answer. you have 52 dimes and 18 nickels.

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhh g's im sorry for being so stupid

OpenStudy (anonymous):

or rather, 52 nickels and 18 dimes.

OpenStudy (anonymous):

im trying really hard to learn this stuff but it look slike chinese to me (ya know a foreign language)

OpenStudy (anonymous):

right. you almost did it correctly. but you have to be careful while subtracting and adding. if you miss that you multiplied something by 1 rather than 0.1, it could lead to big problems down the road.

OpenStudy (anonymous):

thanks for the tip....i really really appreciate your help

OpenStudy (anonymous):

also, it helps if you put yourself in the position of the problem. suppose YOU had x amount of dimes and Y amount of nickels....... and so on. it helps you keep track of what you are trying to find, and not end up with nonsensical answers like -958. because you physically cannot have negative 900 odd nickels.

OpenStudy (anonymous):

you are welcome. all the best!

OpenStudy (anonymous):

thank you!